Heinz School Carnegie Mellon University

Intro to Econometric Theory, 90-906 Spring, 2005-6

Syllabus

Lectures  M 5:30-8:20, HBH 1002
Instructor  Bill Vogt

HBH 2116D, 412.268.1843 wilibear@andrew.cmu.edu

Office Hours M 3:00-5:00
TA  Jian Li

HBH 243, 412.268.1415 jianl@andrew.cmu.edu

TA Ofc Hours TR 10:30-12:00
Textbook  Greene, William H (2000) Econometric Analysis,

5^{th} edition. Upper Saddle River, NJ: Prentice Hall.

(optional) Kennedy, P. (1998) A Guide to Econometrics. Cambridge, MA MIT Press.

Website  http://www.andrew.cmu.edu/course/90-906/index.htm

1 Course Objectives

Our objective is to build a thorough understanding of multivariate linear regression and (time permitting) instrumental variables and discrete dependent variables models.

2 Prerequisites

You are expected to know elementary probability theory, sampling theory, linear algebra, calculus, and bivariate regression. We will briefly review some linear algebra and sampling theory at the beginning of the course, and the book has an excellent review in chapters 1-4. The formal prerequisite is 90-905, an introductory Ph.D. level statistics course, or the consent of the instructor.

3 Topics

The topic for the course will be the linear regression model. Roughly, we will cover Greene, chapters 1-12 and appendix A.

4 Grading

The grade will be determined by performance on homework assignments (30%), a midterm (30%), and the final exam (40%). You are primarily responsible for the content of the lectures, including any handouts. You are also responsible for the relevant sections of the book — most likely chapters 1-9,11-13. There are likely to be differences in both content and notation between the lectures and the book.

5 Various Class Policies

5.1 Policy on Cooperation

• You may cooperate as much as you like on homework assignments; however, each individual must submit a separate assignment, and each individual will be separately graded on that assignment.

• You may not cooperate while taking exams.

5.2 Policy on Aids During Exams

• All exams are open book. You may use text, notes, calculators, computers, reference materials, etc.

• You may not communicate or cooperate with anyone on the exam.

• Please do not use the open book policy as a substitute for studying. If you need to look in your book, notes, etc for instructions on HOW TO DO the exam, you will surely run out of time.

5.3 Policy on Format of Assigned Work

• All work must be legible. Illegible is the equivalent of completely wrong.

• Spelling, grammar, style, etc do not “count” per se. However, anything which I (or the TAs) cannot understand is wrong. Poor spelling, grammar, style, etc are often confusing.

5.4 Policy on Lateness

• Due and return dates for assignments are in the syllabus in the schedule section. Unless the instructor or TA announce otherwise, these dates are binding.

• Late assignments are not accepted without either prior arrangement or compelling reason.

• Assignments and homeworks may be picked up from the TA. You are responsible for picking them up. You are responsible for keeping them to compare against your final grade. I will use whatever grade I have recorded for your work (even if you think there is an error) unless you can show me the graded work so that I can see the error.

• You are responsible for collecting your graded work. Claims by students similar to “I turned in homework #2, but I don’t have a copy of it because you did not return it to me!” must be made within one week of the return date of the assignment. If you fail to make the complaint, I will use whatever grade I have in my records.